Find at least one solution to the following equation: sin(x^2 −1)/ 1−sin(x^2 −1) = sin(x) + sin^2(x) + sin^3(x) + sin^4(x) +···
1
Expert's answer
2020-11-17T06:29:30-0500

sin(x2-1)/(1-sin(x2-1))=sin(x)+sin2(x)+sin3(x)+sin4(x)+...

sin(x2-1)/(1-sin(x2-1))=sin(x)/(1-sin(x))

sin(x2-1)(1-sin(x))=sin(x)(1-sin(x2-1))

sin(x2-1)-sin(x2-1)sin(x)=sin(x)-sin(x2-1)sin(x)

sin(x2-1)-sin(x)--sin(x2-1)sin(x)+sin(x2-1)sin(x)=0

sin(x2-1)-sin(x)=0

2sin((x2-1-x)/2)cos((x2-1+x)/2)=0

sin((x2-1-x)/2)=0

(x2-1-x)/2=0

x2-x-1=0

x=(1±√5)/2